OpenStudy (anonymous):

(3x-2)^2 again ive got the 9x^2+4 how do i get the -12x

OpenStudy (anonymous):

(3x-2)^2 =(3x-2)(3x-2) =3x(3x-2)-2(3x-2) =9x^2-6x-6x+4 =9x^2-12x+4

OpenStudy (anonymous):

ok i understand this one thanks